# Quantum Spin Liquids in Pyrochlore Magnets With Non-Kramers Local Moments

### v1

@@ Appendix C Constraints on the Mean-Field Amplitudes
where the $O(4)$ matrices $\mathcal{R}_{\mathcal{O}}^{\gamma}$ are determined by $U_{\mathcal{O}}$ and the phase factors in Eqs. (40a-40b). This shows that it is enough to specify $u^{\gamma}_{ij}$ on a single bond $(i_{0},j_{0})$ to determine all the remaining $u^{\gamma}_{ij}$.Finally, since we have specialized to $\mathbb{Z}_{2}$ spin liquids, the phases are either $1$ or $-1$, so $(\mathcal{R}^{\gamma}_{\mathcal{O}})^{*}=\mathcal{R}^{\gamma}_{\mathcal{O}}$.
@@
The last set of constraints arises because a bond can be mapped to itself in several inequivalent ways [64]. First, we note that the hopping and pairing mean-field amplitudes satisfy the following symmetry properties
@@
which follows from the definitions (12a-12d). A group theory analysis [64] shows that
@@ Appendix D Mean-Field Decoupling Scheme
We then choose the values of $A,B,C^{x},C^{y},C^{z},N$ so that the decoupling matrix $\xi_{ij}$ is always positive-definite, meaning these values depend on the signs of $J_{zz}$ and $J_{\pm}$. This is done so that the mean field Hamiltonian is bounded below in the amplitudes $\chi,E^{x},\ldots,D^{z}$, and also to ensure $\xi_{ij}$ is always invertible. The boundedness criterion is not strictly necessary, but it allows one to use numerical methods such as simulated annealing to solve the self-consistent equations. However, it is important that $\xi_{ij}$ is invertible for the generalized self-consistent equations (appendix E) to make sense.
@@
$[L^{h}_{ij}]_{1,1}&=|J_{\pm}|\\ [L^{h}_{ij}]_{2,2}=[L^{h}_{ij}]_{3,3}&=\tfrac{1}{4}|J_{zz}|+|J_{\pm}|+\left(1+\tfrac{1}{4}\textrm{Im}\gamma_{ij}\right)|J_{\pm\pm}|\\ [L^{h}_{ij}]_{2,3}=[L^{h}_{ij}]_{3,2}&=\left(1+\tfrac{1}{4}\textrm{Im}\gamma_{ij}\right)|J_{\pm\pm}|\\ [L^{h}_{ij}]_{4,4}&=\tfrac{1}{4}|J_{zz}|+\tfrac{3}{2}|J_{\pm}|+\tfrac{1}{2}|J_{\pm\pm}|\\ [L^{p}_{ij}]_{1,1}&=|J_{\pm}|+|J_{\pm\pm}|\\ [L^{p}_{ij}]_{2,2}=[L^{p}_{ij}]_{3,3}&=\tfrac{1}{2}|J_{\pm}|+\left(2+\tfrac{1}{2}\textrm{Re}\gamma_{ij}+\tfrac{1}{2}\textrm{Im}\gamma_{ij}\right)|J_{\pm\pm}|\\ [L^{p}_{ij}]_{2,3}=[L^{p}_{ij}]_{3,2}&=\left(1+\tfrac{1}{4}\textrm{Im}\gamma_{ij}\right)|J_{\pm\pm}|\\ [L^{p}_{ij}]_{4,4}&=\tfrac{1}{4}|J_{zz}|+\tfrac{1}{2}|J_{\pm}|+\tfrac{3}{2}|J_{\pm\pm}|\end{aligned}$ $[L^{h}_{ij}]_{1,1}&=\tfrac{1}{4}|J_{zz}|\\ [L^{h}_{ij}]_{2,2}=[L^{h}_{ij}]_{3,3}&=\tfrac{1}{4}|J_{zz}|+\left(1+\tfrac{1}{4}\textrm{Im}\gamma_{ij}\right)|J_{\pm\pm}|\\ [L^{h}_{ij}]_{2,3}=[L^{h}_{ij}]_{3,2}&=\left(1+\tfrac{1}{4}\textrm{Im}\gamma_{ij}\right)|J_{\pm\pm}|\\ [L^{h}_{ij}]_{4,4}&=\tfrac{1}{4}|J_{zz}|+\tfrac{1}{2}|J_{\pm}|+\tfrac{1}{2}|J_{\pm\pm}|\\ [L^{p}_{ij}]_{1,1}&=|J_{\pm}|+|J_{\pm\pm}|\\ [L^{p}_{ij}]_{2,2}&=\tfrac{1}{4}|J_{zz}|+\tfrac{1}{2}|J_{\pm}|+\left(2+\tfrac{1}{2}\textrm{Re}\gamma_{ij}+\tfrac{1}{4}\textrm{Im}\gamma_{ij}\right)|J_{\pm\pm}|\\ [L^{p}_{ij}]_{3,3}&=\tfrac{1}{4}|J_{zz}|+\tfrac{1}{2}|J_{\pm}|+\left(2-\tfrac{1}{2}\textrm{Re}\gamma_{ij}+\tfrac{1}{4}\textrm{Im}\gamma_{ij}\right)|J_{\pm\pm}|\\ [L^{p}_{ij}]_{2,3}=[L^{p}_{ij}]_{3,2}&=\left(1-\tfrac{1}{4}\textrm{Im}\gamma_{ij}\right)|J_{\pm\pm}|\\ [L^{p}_{ij}]_{4,4}&=\tfrac{1}{2}|J_{\pm}|+\tfrac{3}{2}|J_{\pm\pm}|\end{aligned}$
@@ E.1 The Standard Approach and Its Issues
For brevity, we have denoted the expectation value of the vector $\bm{B}$ by $\bm{X}$ (X for eXpectation value). Then,
@@
provided $\xi_{k\ell}$ is invertible. We call Eq. (56b) the “naive self-consistent equations”, which holds for all bonds $(k,\ell)$. In the notation of previous work such as [64], Eq. (56b) reads
@@
Now, we would like to impose PSG constraints on the ansatz. As argued in the main text, this can modify the form of the self-consistent equations beacuse the symmetries of the ansatz mean that the amplitudes on each bond are no longer independent. If the ansatz respects the PSG, there are only a few independent mean-field amplitudes (say, the set of $A_{01}$ for all $\hat{A}$), and the remaining $A_{ij}$ can depend on them in nontrivial ways. Therefore, one should apply the PSG constraints to the ansatz first and then minimize the energy with respect to the independent mean-field amplitudes. It can happen that the two approaches are equivalent if the decoupling and PSG constraints are simple enough, as was the case in the previously studied dipolar-octupolar model [64].
@@
and we see that Eqs. (59a) and (59b) are not the same as (60). This is to be expected because the first approach does not actually work in general, as the above example shows. However, if the constraint was instead $g(x,y)=y-x$, then the first approach still works. This is only because we got lucky with the specific constraint. Thus, we are led to suspect that the self-consistent equations are not always (56b).
@@ E.4 The Generalized Self-Consistent Equations
00 0 $=\frac{\partial\langle H_{\textrm{MF}}\rangle}{\partial\bm{{\mathcal{A}}}_{01}^{\dagger}}$ (73a)
@@ F.1 Fourier Transformation of the Mean-Field Hamiltonian
where the $8\times 8$ blocks $H_{h}$ and $H_{p}$ satisfy $H_{h}({{\bm{k}}})=H^{\dagger}_{h}({{\bm{k}}})$ and $H_{p}({{\bm{k}}})=H^{\intercal}(-{{\bm{k}}})$. Therefore, it is enough to specify the upper triangular part of the $H_{\gamma}$ matrices. Explicitly,
@@ F.2 Bogoliubov Transformation
We determine $P({{\bm{k}}})$ by Colpa’s algorithm, with the additional step of reordering the eigenvalues to appear as in Eq. (95b) [75].
@@ F.3 Evaluating Ground State Expectation Values
due to the structure of the matrix $\mathcal{H}({{\bm{k}}})$ [60]. We can then write the Bogoliubov transformation (94) as
@@
[…]
@@
×$\times[\sigma]^{a}_{\alpha\beta}e^{-i{{\bm{k}}}\cdot({{\bm{r}}}-{{\bm{r}}}^{\prime})}\Big{)}$ (104a)
@@
×$\times[i\sigma^{y}\sigma]^{a}_{\alpha\beta}e^{i{{\bm{k}}}\cdot({{\bm{r}}}-{{\bm{r}}}^{\prime})}\Big{)}.$ (104b)

### v2

@@ Appendix C Constraints on the Mean-Field Amplitudes
where the $O(4)$ matrices $\mathcal{R}_{\mathcal{O}}^{\gamma}$ are determined by $U_{\mathcal{O}}$ and the phase factors in Eqs. (40a-40b). This shows that it is enough to specify $u^{\gamma}_{ij}$ on a single bond $(i_{0},j_{0})$ to determine all the remaining $u^{\gamma}_{ij}$.Finally, since we have specialized to $\mathbb{Z}_{2}$ spin liquids, the phases are either $1$ or 1-1, so $(\mathcal{R}^{\gamma}_{\mathcal{O}})^{*}=\mathcal{R}^{\gamma}_{\mathcal{O}}$.
@@
The last set of constraints arises because a bond can be mapped to itself in several inequivalent ways [Desrochers2022]. First, we note that the hopping and pairing mean-field amplitudes satisfy the following symmetry properties
@@
which follows from the definitions (12a-12d). A group theory analysis [Desrochers2022] shows that
@@ Appendix D Mean-Field Decoupling Scheme
We then choose the values of $A,B,C^{x},C^{y},C^{z},N$ so that the decoupling matrix $\xi_{ij}$ is always positive-definite, meaning these values depend on the signs of $J_{zz}$ and $J_{\pm}$. This is done so that the mean-field Hamiltonian is bounded below in the amplitudes $\chi,E^{x},\ldots,D^{z}$, and also to ensure $\xi_{ij}$ is always invertible. The boundedness criterion is not strictly necessary, but it allows one to use numerical methods such as simulated annealing to solve the self-consistent equations. However, it is important that $\xi_{ij}$ is invertible for the generalized self-consistent equations (appendix E) to make sense.
@@
$\begin{aligned} [L^{h}_{ij}]_{1,1}&=|J_{\pm}|\\ [L^{h}_{ij}]_{2,2}=[L^{h}_{ij}]_{3,3}&=\tfrac{1}{4}|J_{zz}|+|J_{\pm}|+\left(1+\tfrac{1}{4}\textrm{Im}\gamma_{ij}\right)|J_{\pm\pm}|\\ [L^{h}_{ij}]_{2,3}=[L^{h}_{ij}]_{3,2}&=\left(1+\tfrac{1}{4}\textrm{Im}\gamma_{ij}\right)|J_{\pm\pm}|\\ [L^{h}_{ij}]_{4,4}&=\tfrac{1}{4}|J_{zz}|+\tfrac{3}{2}|J_{\pm}|+\tfrac{1}{2}|J_{\pm\pm}|\\ [L^{p}_{ij}]_{1,1}&=|J_{\pm}|+|J_{\pm\pm}|\\ [L^{p}_{ij}]_{2,2}=[L^{p}_{ij}]_{3,3}&=\tfrac{1}{2}|J_{\pm}|+\left(2+\tfrac{1}{2}\textrm{Re}\gamma_{ij}+\tfrac{1}{2}\textrm{Im}\gamma_{ij}\right)|J_{\pm\pm}|\\ [L^{p}_{ij}]_{2,3}=[L^{p}_{ij}]_{3,2}&=\left(1+\tfrac{1}{4}\textrm{Im}\gamma_{ij}\right)|J_{\pm\pm}|\\ [L^{p}_{ij}]_{4,4}&=\tfrac{1}{4}|J_{zz}|+\tfrac{1}{2}|J_{\pm}|+\tfrac{3}{2}|J_{\pm\pm}|\end{aligned}$ $\begin{aligned} [L^{h}_{ij}]_{1,1}&=\tfrac{1}{4}|J_{zz}|\\ [L^{h}_{ij}]_{2,2}=[L^{h}_{ij}]_{3,3}&=\tfrac{1}{4}|J_{zz}|+\left(1+\tfrac{1}{4}\textrm{Im}\gamma_{ij}\right)|J_{\pm\pm}|\\ [L^{h}_{ij}]_{2,3}=[L^{h}_{ij}]_{3,2}&=\left(1+\tfrac{1}{4}\textrm{Im}\gamma_{ij}\right)|J_{\pm\pm}|\\ [L^{h}_{ij}]_{4,4}&=\tfrac{1}{4}|J_{zz}|+\tfrac{1}{2}|J_{\pm}|+\tfrac{1}{2}|J_{\pm\pm}|\\ [L^{p}_{ij}]_{1,1}&=|J_{\pm}|+|J_{\pm\pm}|\\ [L^{p}_{ij}]_{2,2}&=\tfrac{1}{4}|J_{zz}|+\tfrac{1}{2}|J_{\pm}|+\left(2+\tfrac{1}{2}\textrm{Re}\gamma_{ij}+\tfrac{1}{4}\textrm{Im}\gamma_{ij}\right)|J_{\pm\pm}|\\ [L^{p}_{ij}]_{3,3}&=\tfrac{1}{4}|J_{zz}|+\tfrac{1}{2}|J_{\pm}|+\left(2-\tfrac{1}{2}\textrm{Re}\gamma_{ij}+\tfrac{1}{4}\textrm{Im}\gamma_{ij}\right)|J_{\pm\pm}|\\ [L^{p}_{ij}]_{2,3}=[L^{p}_{ij}]_{3,2}&=\left(1-\tfrac{1}{4}\textrm{Im}\gamma_{ij}\right)|J_{\pm\pm}|\\ [L^{p}_{ij}]_{4,4}&=\tfrac{1}{2}|J_{\pm}|+\tfrac{3}{2}|J_{\pm\pm}|\end{aligned}$
@@ E.1 The Standard Approach and Its Issues
For brevity, we have denoted the expectation value of the vector $\bm{B}$ by $\bm{X}$. Then,
@@
provided $\xi_{k\ell}$ is invertible. We call Eq. (56b) the “naive self-consistent equations”, which holds for all bonds $(k,\ell)$. In the notation of previous work such as [Desrochers2022], Eq. (56b) reads
@@
Now, we would like to impose PSG constraints on the ansatz. As argued in the main text, this can modify the form of the self-consistent equations beacuse the symmetries of the ansatz mean that the amplitudes on each bond are no longer independent. If the ansatz respects the PSG, there are only a few independent mean-field amplitudes (say, the set of $A_{01}$ for all $\hat{A}$), and the remaining $A_{ij}$ can depend on them in nontrivial ways. Therefore, one should apply the PSG constraints to the ansatz first and then minimize the energy with respect to the independent mean-field amplitudes. It can happen that the two approaches are equivalent if the decoupling and PSG constraints are simple enough, as was the case in the previously studied dipolar-octupolar model [Desrochers2022].
@@
and we see that Eqs. (59a) and (59b) are not the same as (60). This is to be expected because the first approach does not actually work in general, as the above example shows. However, if the constraint was instead g(x,y)=yxg(x,y)=y-x, then the first approach still works. This is only because we got lucky with the specific constraint. Thus, we are led to suspect that the self-consistent equations are not always (56b).
@@ E.4 The Generalized Self-Consistent Equations
$0$ $=\frac{\partial\langle H_{\textrm{MF}}\rangle}{\partial\bm{{\mathcal{A}}}_{01}^{\dagger}}$ (73a)
@@ F.1 Fourier Transformation of the Mean-Field Hamiltonian
where the $8\times 8$ blocks $H_{h}$ and $H_{p}$ satisfy $H_{h}({{\bm{k}}})=H^{\dagger}_{h}({{\bm{k}}})$ and $H_{p}({{\bm{k}}})=H_{p}^{\intercal}(-{{\bm{k}}})$. Therefore, it is enough to specify the upper triangular part of the $H_{\gamma}$ matrices. Explicitly,
@@ F.2 Bogoliubov Transformation
We determine $P({{\bm{k}}})$ by Colpa’s algorithm, with the additional step of reordering the eigenvalues to appear as in Eq. (95b) [Colpa1978].
@@ F.3 Evaluating Ground State Expectation Values
due to the structure of the matrix $\mathcal{H}({{\bm{k}}})$ [Liu2019]. We can then write the Bogoliubov transformation (94) as
@@
where the composite sublattice-spin index $(\mu\alpha)$ represents a row or column index ranging from $1$ to $8$. In particular, if the spin index $\alpha$ is mapped to $\uparrow\ \mapsto 1$, $\downarrow\ \mapsto 2$, the composite index $(\mu\alpha)$ is equivalent to the usual matrix index $2\mu+\alpha$. Since the ground state $|0\rangle$ is defined by $\gamma_{{{\bm{k}}},\mu,\alpha}|0\rangle=0$ for all ${{\bm{k}}},\mu,\alpha$, the only $\gamma$ bilinear that does not vanish immediately is
@@
$\times[\sigma]^{a}_{\alpha\beta}e^{-i{{\bm{k}}}\cdot({{\bm{r}}}-{{\bm{r}}}^{\prime})}\Big)$ (104a)
@@
$\times[i\sigma^{y}\sigma^{a}]_{\alpha\beta}e^{i{{\bm{k}}}\cdot({{\bm{r}}}-{{\bm{r}}}^{\prime})}\Big).$ (104b)